OpenStudy (anonymous):

If the parent function is y=1/x , describe the change in the equation y=1/x-7.

OpenStudy (anonymous):

Moves 7 units up Moves 7 units down Moves 7 units to the left Moves 7 units to the right

OpenStudy (anonymous):

I said C

OpenStudy (anonymous):

you really are troublesome now we have to graph it

OpenStudy (anonymous):

hey u there

OpenStudy (anonymous):

yes

OpenStudy (yrelhan4):

y=1/(x-7).. is it this? or (1/x)-7 ?

OpenStudy (anonymous):

http://prntscr.com/1dh3yn

OpenStudy (yrelhan4):

ok you need to graph it.. take any 3 values of x and find the values of y for them..

OpenStudy (yrelhan4):

and then plot it.. do it for y=1/x and y=1/x-7.. once you have graphed it, you'll get to know the answer.

OpenStudy (yrelhan4):

it isnt a.. graph it yourself using 3-4 values of x,y and in that link type 1/(x-7)..

OpenStudy (yrelhan4):

and its C.. you're right @iSuckAtM4th

OpenStudy (yrelhan4):

eh wait. it isnt C..

OpenStudy (anonymous):

hehe

OpenStudy (yrelhan4):

GRAPH IT..

OpenStudy (anonymous):

I'm graphing it now, But I'm getting kinda of C

OpenStudy (anonymous):

i think its d idk

OpenStudy (anonymous):

Nah im an idiot

OpenStudy (yrelhan4):

ok look. for x=7 y=1/0(or undefined or infinity) for x=infinity y=0 (or tends to 0) get it so far?

OpenStudy (yrelhan4):

now you just need to plot the points..